Hours after a governor was indicted on several counts of fraud and extortion, two reputable research firms polled citizens frA poll found that people from the north now have relatively warm feelings toward people from the south. A poll, conducted byUsing the given margin of error, confidence level, and likely range for the observed value of the sample proportion, obtain aFor the specified margin of error, confidence level, and educated guess for the observed value, obtain a sample size that wilFor the specified margin of error and confidence level, obtain a sample size that will ensure a margin of error of at most th
Answer #1

My dear student I am allow to solve the one question at a time. So I solved the first question in the picture.

60-4.8= 55.2 Corred oblim is to 60+4.8 = 64-8 and the Becond confidence interval 55-4.9 - 50-8 to 55+4,2 = 59.2 Overlab